Phyllis invested 27000 dollars, a portion earning a simple interest rate of 4 percent per year and the rest earning a rate of 7 percent per year. After one year the total interest earned on these investments was 1530 dollars. How much money did she invest at each rate?
Answer by stanbon(75887)

Phyllis invested 27000 dollars, a portion earning a simple interest rate of 4 percent per year and the rest earning a rate of 7 percent per year. After one year the total interest earned on these investments was 1530 dollars. How much money did she invest at each rate?
---------------------------
Equation:
Quantity Equation: x + y = 27000
Interest Equation:4x +7y = 153000
-----------------------------------
Modify to eliminate:
4x + 4y = 4*27000
4x + 7y = 153000
-------
3y = 45000
y = $15000 (amt invested at 4%
x = 27000-15000 (amt. invested at 7%
----------------------------------------
Cheers,
Stan H.